John likes dark coffe beans, which cost $8.50/kg. Brad likes mild coffe beans which cost $9.20/kg. They buy a 5.5kg made from a blend of both type of beans that cost $48.50. How many kg of each type of bean went into the mixture?
Answer by checkley77(12844)

9.20X+8.50(5.5-X)=48.50
9.20X+46.75-8.50X=48.50
.70X=48.50-46.75
.70X=1.75
X=1.75/.70
X=2.5 LBS. OF $9.20 COFFEE.
5.5-.5=3 LBS OF $8.50 COFFEE.
PROOF:
9.20*2.5+8.50*3=48.50
23+25.5=48.50
48.50=48.50